# Non-Convex Over-the-Air Heterogeneous Federated Learning: A Bias–Variance Trade-off

### v3

@@ Abstract
Over-the-air (OTA) federated learning (FL) has been well recognized as a scalable paradigm that exploits the waveform superposition of the wireless multiple-access channel to aggregate model updates simultaneously. Existing OTA-FL designs largely enforce zero-bias model updates by either assuming homogeneous wireless conditions (equal path loss across devices) or forcing zero bias updates to guarantee convergence. Under heterogeneous wireless scenarios, however, such designs are constrained by the weakest device and degrade the update variance. Moreover, prior analyses of biased OTA-FL largely address convex objectives, while most modern AI models are highly non-convex. Motivated by these gaps, we study OTA-FL with stochastic gradient descent (SGD) for smooth non-convex objectives under wireless heterogeneity. We develop novel OTA-FL SGD updates that allow a structured, time-invariant model bias while facilitating reduced variance updates. We derive a finite-time stationarity bound (expected time average squared gradient norm) that explicitly reveals a bias–variance trade-off. To optimize this trade-off, we pose a non-convex joint OTA power-control design and develop an efficient successive convex approximation (SCA) algorithm that requires only statistical CSI at the base station. Experiments on a non-convex image classification task validate the approach: the SCA-based design accelerates convergence via an optimized bias and improves generalization over prior OTA-FL baselines.
@@ I Introduction
Federated Learning (FL) has emerged as a key paradigm for privacy-preserving distributed learning, allowing a massive number of devices to collaboratively train a model without sharing raw data. In a standard FL setup, $N$ clients with private datasets coordinate with a central parameter server (PS), e.g., an edge or cloud node, by transmitting only model parameters or gradients [FL_Iot_Survey, FL_survey]. The objective is to solve
$\mathbf{w}^{*}\;=\;\arg\min_{\mathbf{w}\in\mathbb{R}^{d}}F(\mathbf{w})\;\triangleq\;\frac{1}{N}\sum_{m\in[N]}f_{m}(\mathbf{w}),$ (P)
@@
where $f_{m}(\mathbf{w})$ represents the local objective function of device $m$, and $F(\mathbf{w})$ is the global objective (loss) function. Problem (P) is typically solved by first-order methods (e.g., mini-batch SGD): in each round, the PS broadcasts the current model $\mathbf{w}_{t}$, devices compute local gradients on their data and transmit them wirelessly to the PS. The PS then aggregates the received updates to obtain $\mathbf{w}_{t+1}$ and re-broadcasts it to begin the next round. This process repeats for multiple rounds until the global objective converges [Fedavg].
@@
Nevertheless, realizing FL over wireless networks requires iterative exchange of high-dimensional gradients over noisy, bandwidth-constrained wireless channels, and hence causes a major performance bottleneck. A growing body of work, therefore, studies FL over wireless channels with robustness to fading and noise [OTA_FL, FL_fading, One_bit_FL, Biased_OTA_FL_ICC], and proposes device-scheduling strategies that account for channel gain fluctuations [Sched_policies, ADFL, BB_FL]. Among these, over-the-air (OTA) computation has emerged as a particularly scalable approach for wireless FL. OTA-FL works on the principle of waveform superposition on the multiple-access channel (MAC), whereby simultaneous transmissions yield “single-shot” model or gradient aggregation at the PS, enabling fast FL updates [OTA_FL, FL_fading, BB_FL]. To ensure unbiased estimation via OTA aggregation, the typical requirement is to align phases and equalize scales via device pre-scalers and a PS post-scaler. Over fading MACs, this coherent aggregation effectively requires channel inversion at the devices, and hence the feasible pre-scaling is dictated by the “weakest” device.
@@
Thresholding and related heuristics have been proposed to mitigate this issue in OTA-FL literature [FL_fading, BB_FL]. However, most prior OTA-FL analyses either assume wireless homogeneity (equal average path loss across devices) to ensure zero average bias or explicitly enforce zero bias to facilitate convergence guarantees [FL_fading, OTA_FL, One_bit_FL, OTA_FL_H_data]. In realistic heterogeneous deployments, where devices experience varying large-scale channel conditions, imposing zero bias can severely increase model update variance (due to weak devices), whereas relaxing the zero bias design constraint, as done in [Opt_Power_Control_OTA_Comp, Opt_power_control_OTA_FL, OTA_FL_Optimization], can introduce unpredictable bias and weak convergence guarantees. In addition, much of the existing convergence theory for OTA-FL is developed under (strongly) convex objectives, see e.g.,[OTA_FL_H_data, Opt_power_control_OTA_FL, Biased_OTA_FL_ICC], which is inconsistent with the modern non-convex ML models deployed in practice.
@@
To fill these gaps, we study OTA-FL under wireless heterogeneity with SGD updates for general smooth non-convex objectives. Building on and extending our prior work on biased OTA-FL in the convex setting [Biased_OTA_FL_ICC], we design OTA-FL updates that admit a structured, time-invariant model bias to reduce update variance, and we generalize the analysis to non-convex objectives. We establish a finite-time stationarity convergence bound (expected average squared gradient norm) that reveals a bias-variance trade-off. Building on this bound, we pose a joint OTA power-control problem and, unlike the heuristic designs proposed in [Biased_OTA_FL_ICC], develop an efficient successive convex approximation (SCA) algorithm that operates with statistical channel state information (CSI) at the PS, avoiding the heavy overhead of global instantaneous CSI. Finally, experiments on a non-convex learning task corroborate the theory, showing that the SCA-based design accelerates convergence by optimizing the allowed bias and achieves a target generalization performance faster than prior OTA-FL baselines.
@@ II System Model and over-the-air FL
We consider a wireless network of $N$ distributed devices and a base station (also acting as the PS) aiming to collaboratively learn an FL model, as shown in Fig. 1. Each device $m$ owns a private local dataset $\mathcal{D}_{m}$ containing $D$ data points, with associated local objective function $f_{m}(\mathbf{w})=\frac{1}{D}\,\sum_{\boldsymbol{\xi}\in\mathcal{D}_{m}}\phi(\mathbf{w},\boldsymbol{\xi})$, only computable at device $m$, where $\phi(\mathbf{w},\cdot)$ is the sample-wise loss function, $\boldsymbol{\xi}$ is a data point and $\mathbf{w}\in\mathbb{R}^{d}$ is the $d$-dimensional learning parameter. The goal is to solve (P) via distributed SGD over multiple FL rounds. Specifically, at the start of round $t$, the PS broadcasts the latest FL model $\mathbf{w}_{t}$ to each device in the network. Each device $m$ then uses a randomly drawn mini-batch $\mathcal{B}_{m,t}\subseteq\mathcal{D}_{m}$ of fixed size $|\mathcal{B}_{m,t}|=B$ to estimate its local gradient $\boldsymbol{g}_{m,t}=\frac{1}{B}\sum_{\boldsymbol{\xi}\in\mathcal{B}_{m,t}}\phi(\mathbf{w}_{t},\boldsymbol{\xi})$. Subsequently, each device uploads its estimated local gradient to the PS. Ideally, the PS aims to aggregate these gradients without error to compute the true global stochastic gradient:
@@
$\mathbf{w}_{t+1}=\mathbf{w}_{t}-{\eta}\overline{\boldsymbol{g}}_{t}.$ (2)
This procedure is repeated until a target accuracy is reached or a fixed number of FL rounds is completed. Nonetheless, computing the global gradient in (1) requires noiseless aggregation of all the local gradients, i.e., they need to be perfectly received at the PS, and aggregated with uniform weights $\frac{1}{N}$. However, in practice, local gradients are obtained through a noisy wireless channel, introducing errors in the estimation of the global gradient $\overline{\boldsymbol{g}}_{t}$, which is discussed next.
@@ II-A Over-the-air transmission over a fading MAC
We model the channel from device $m$ to the PS in round $t$ as flat Rayleigh fading, $h_{m,t}\sim\mathcal{CN}(0,\Lambda_{m})$, independent across devices and i.i.d. over $t$. Here, the parameter $\Lambda_{m}$ represents the average channel gain, which is dependent on large-scale propagation conditions and is assumed to be constant throughout the FL runtime. We assume the PS has knowledge of this statistical CSI, i.e., the set $\{\Lambda_{m}\}$, but not the instantaneous CSI $\{h_{m,t}\}$. Critically, unlike existing works [FL_fading, OTA_FL, One_bit_FL, OTA_FL_H_data] assuming identical average path loss across devices ($\Lambda_{m}{=}\Lambda_{n},\forall m,n\in[N]$}), we consider a heterogeneous wireless environment where devices experience distinct path losses. We consider local gradient uplink transmission using the OTA aggregation scheme. OTA-FL exploits the superposition property of the wireless multiple-access channel to perform joint computation-and-communication, enabling single-shot aggregation at the PS. Following [OTA_FL, FL_fading, BB_FL, One_bit_FL, NCOTA, OTA_FL_H_data], we assume devices are perfectly synchronized during gradient upload. Denoting by $\mathbf{x}_{m,t}$ the signal transmitted by device $m$ in round $t$, the PS receives the signal
@@
Here, $E_{s}$ is the maximum per-sample energy budget, and $G_{\text{max}}$ is an upper bound on $\|\boldsymbol{g}_{m,t}\|$ (see Assumption 2). By construction, a device remains silent when $|h_{m,t}|<\frac{G_{\text{max}}\gamma_{m}}{\sqrt{dE_{s}}}$ to ensure the energy constraint $\|\mathbf{x}_{m,t}\|^{2}/d\leq E_{s}\,,\forall m,t$, where $d$ is the model dimension. The decision in (5) is decentralized using the instantaneous uplink CSI $h_{m,t}$, which can be acquired with negligible overhead via a downlink pilot, by leveraging channel reciprocity [FL_d2d]. With this design, the PS estimates the global gradient (1) as
@@ II-B Biased Over-the-Air-FL
in place of the ideal SGD update in (2). Conditioned on $\mathbf{w}_{t}$ and the mini-batch selections, the received signal satisfies $\mathbb{E}[\mathbf{y}_{t}]=\sum_{m\in[N]}\alpha_{m}\,\boldsymbol{g}_{m,t}$ with $\alpha_{m}\triangleq\gamma_{m}\,\mathbb{E}[\chi_{m,t}]={\gamma_{m}\exp\{-\tfrac{\gamma_{m}^{2}G_{\text{max}}^{2}}{d\,\Lambda_{m}\,E_{s}}\}}$, where the expectation is over channel fading and receiver noise. Choosing the post-scaler $\alpha\triangleq\sum_{m\in[N]}\alpha_{m}$ yields a global gradient estimate such that the expected gradient estimate is a convex combination of local gradients,
$\tilde{\boldsymbol{g}}_{t}\;\triangleq\;\mathbb{E}[\hat{\boldsymbol{g}}_{t}\mid\mathbf{w}_{t}]\;=\;\sum_{m\in[N]}p_{m}\,\boldsymbol{g}_{m,t},$ (8)
with weights $p_{m}\triangleq\alpha_{m}/\alpha$, $0\leq p_{m}\leq 1$, $\sum_{m}p_{m}=1$. We interpret $p_{m}$ as the average participation level of device $m$ induced by the OTA-FL design. Consequently, the update rule in (7) acts as an SGD method that, on average, descends along $\tilde{\boldsymbol{g}}_{t}$ in place of the ideal aggregate $\overline{\boldsymbol{g}}_{t}$ from (1). Therefore, these updates minimize a different objective function than (P), on average, given by
@@
It is to be highlighted that prior OTA-FL works [Dig_vs_Analog, ADFL, FL_fading, OTA_FL, One_bit_FL, OTA_FL_H_data] either assume homogeneous large-scale wireless conditions or enforce zero-bias aggregation, thereby ensuring uniform participation $p_{m}=\frac{1}{N}$ for all $m\in[N]$, so that minimizing ($\tilde{\text{P}}$) becomes equivalent to (P). While effective under homogeneous conditions, these designs become constrained by the weakest channel device in heterogeneous wireless networks, leading to high-variance model updates (see, e.g., [BB_FL, Biased_OTA_FL_ICC, Opt_power_control_OTA_FL]). Moreover, under wireless heterogeneity, such schemes effectively minimize a different objective, leading to objective inconsistency [Obj_incons] and a non-negligible model bias. As a result, their convergence guarantees cannot be applied to the scenario considered in this paper.
Our proposed OTA-FL aggregation design generalizes existing schemes by introducing a controllable, non-zero average bias, thereby subsuming zero-bias designs ($p_{m}=\tfrac{1}{N}$) as a special case. This formulation exposes an explicit bias–variance trade-off that can be optimized to improve convergence. While some recent works [Opt_power_control_OTA_FL, OTA_FL_Optimization] also consider biased OTA-FL, they typically allow a generic, unstructured bias that is difficult to analyze and control. In contrast, building on our prior work [Biased_OTA_FL_ICC], we employ a structured, time-invariant bias that admits tractable finite-time guarantees. Note that, unlike the prior study, which treated (strongly) convex objectives and offered heuristic designs, here we derive a convergence bound for non-convex objectives and optimize the bias-variance trade-off via an efficient SCA method, which is done in the next section.

### v4

@@ Abstract
Over-the-air (OTA) federated learning (FL) has been well recognized as a scalable paradigm that exploits the waveform superposition of the wireless multiple-access channel to aggregate model updates simultaneously. Existing OTA-FL designs largely enforce zero-bias model updates by either assuming homogeneous wireless conditions (equal path loss across devices) or forcing zero bias updates to guarantee convergence. Under heterogeneous wireless scenarios, however, such unbiased designs are constrained by the worst-channel device and suffer from high variance in the updates. Moreover, prior analyses of biased OTA-FL largely focus on convex objectives, whereas most modern machine-learning models are highly non-convex. Motivated by these gaps, we study OTA-FL with stochastic gradient descent (SGD) for smooth non-convex objectives under wireless heterogeneity. We develop novel OTA-FL SGD updates that allow a structured, time-invariant model bias while facilitating reduced variance in the updates. We also derive a finite-time stationarity bound (expected time average squared gradient norm) that explicitly reveals a bias–variance trade-off. To optimize this trade-off, we pose a non-convex joint OTA power-control design and develop an efficient successive convex approximation (SCA) algorithm that requires only statistical CSI of the devices at the base station. Experiments on a non-convex image classification task validate the approach: the SCA-based design accelerates convergence via an optimized bias and improves generalization over prior OTA-FL baselines.
@@ I Introduction
Federated Learning (FL) has emerged as a key paradigm for privacy-preserving distributed learning, allowing a massive number of devices to collaboratively train a model without sharing raw data. In a standard FL setup, $N$ clients with private datasets coordinate with a central parameter server (PS), e.g., an edge or cloud server, by transmitting only model parameters or gradients [15, 10]. The objective is to solve
$\mathbf{w}^{*}=\arg\min_{\mathbf{w}\in\mathbb{R}^{d}}F(\mathbf{w}),\text{where }F(\mathbf{w})\triangleq\frac{1}{N}\sum_{m\in[N]}f_{m}(\mathbf{w}),$ (P)
@@
$f_{m}(\mathbf{w})$ represents the local objective function of device $m$, and $F(\mathbf{w})$ is the global objective (loss) function. Problem (P) is typically solved by first-order optimization methods (e.g., mini-batch SGD): in each round, the PS broadcasts the current model $\mathbf{w}_{t}$, devices compute local gradients on their data and transmit them wirelessly to the PS. The PS then aggregates the received updates to obtain $\mathbf{w}_{t+1}$ and re-broadcasts it to begin the next round. This process repeats for multiple rounds until the global objective converges [13].
@@
Nevertheless, realizing FL over wireless networks requires iterative exchange of high-dimensional parameters over noisy, bandwidth-constrained wireless channels, thereby posing a major performance bottleneck. A growing body of work, therefore, studies FL over wireless channels with robustness to fading and noise [21, 3, 23, 1], and proposes device-scheduling strategies that account for channel gain fluctuations [20, 17, 24]. Among these approaches, over-the-air (OTA) computation has emerged as a particularly scalable solution for wireless FL. OTA-FL works on the principle of waveform superposition on the multiple-access channel (MAC), whereby simultaneous transmissions yield “single-shot” parameter aggregation at the PS, enabling fast FL updates [21, 3, 24]. To ensure unbiased estimation via OTA aggregation, the typical requirement is to align phases and equalize scales via device pre-scalers and a PS post-scaler. Over fading MACs, this coherent aggregation effectively requires channel inversion at the devices, and hence the feasible pre-scaling is dictated by the “weakest” device.
@@
Thresholding and related heuristics have been proposed to mitigate this issue in OTA-FL literature [3, 24]. However, most prior OTA-FL analyses either assume wireless homogeneity (equal average path loss across devices) to ensure zero average bias FL updates or explicitly enforce zero bias to facilitate convergence guarantees [3, 21, 23, 16]. In realistic heterogeneous deployments, where devices experience varying large-scale channel conditions, imposing zero bias can substantially increase model update variance (due to weak-channel devices). Conversely, relaxing the zero-bias design constraint, as in [5, 6, 25], can introduce unstructured bias that is difficult to quantify, resulting in weaker convergence guarantees. In addition, much of the existing convergence theory for OTA-FL is developed under (strongly) convex objectives, see e.g.,[16, 6, 1], which is inconsistent with the non-convex ML models deployed in modern practice.
@@
To fill these gaps, we study OTA-FL under wireless heterogeneity with SGD updates for general smooth non-convex objectives. Building on and extending our prior work on biased OTA-FL in the strongly convex setting [1], we design OTA-FL updates that admit a structured, time-invariant bias to reduce the variance of the updates, and we generalize the analysis to non-convex objectives. We establish a finite-time stationarity convergence bound (expected average squared gradient norm) that reveals a bias-variance trade-off. Building on this bound, we pose a joint OTA power-control problem and, unlike the heuristic designs proposed in [1], develop an efficient successive convex approximation (SCA) algorithm that requires only statistical channel state information (CSI) at the PS, avoiding the heavy overhead of global instantaneous CSI. Finally, experiments on a non-convex learning task corroborate the theory, showing that the SCA-based design accelerates convergence by optimizing the allowed bias and achieves a target generalization performance faster than prior OTA-FL baselines.
@@ II System Model and over-the-air FL
We consider a wireless network of $N$ distributed devices and a base station (also acting as the PS) aiming to collaboratively learn an FL model, as shown in Fig. 1. Each device $m$ owns a private local dataset $\mathcal{D}_{m}$ containing $D$ data points, with associated local objective function $f_{m}(\mathbf{w})=\frac{1}{D}\,\sum_{\boldsymbol{\xi}\in\mathcal{D}_{m}}\phi(\mathbf{w},\boldsymbol{\xi})$, only computable at device $m$, where $\phi(\mathbf{w},\cdot)$ is the sample-wise loss function, $\boldsymbol{\xi}$ is a data point and $\mathbf{w}\in\mathbb{R}^{d}$ is the $d$-dimensional learning parameter. The goal is to solve (P) via distributed SGD over multiple FL rounds. Specifically, at the start of round $t$, the PS broadcasts the latest FL model $\mathbf{w}_{t}$ to each device in the network. Each device $m$ then uses a randomly drawn mini-batch $\mathcal{B}_{m,t}\subseteq\mathcal{D}_{m}$ of fixed size $|\mathcal{B}_{m,t}|=B$ to estimate its local gradient $\boldsymbol{g}_{m,t}=\frac{1}{B}\sum_{\boldsymbol{\xi}\in\mathcal{B}_{m,t}}\nabla\phi(\mathbf{w}_{t},\boldsymbol{\xi})$. Subsequently, each device uploads its estimated local gradient to the PS. Ideally, the PS aims to aggregate these gradients without error to compute the true global stochastic gradient:
@@
$\mathbf{w}_{t+1}=\mathbf{w}_{t}-{\eta}\overline{\boldsymbol{g}}_{t},\quad t\geq 0.$ (2)
This procedure is repeated until a target accuracy is reached or a fixed number of FL rounds is completed. Nonetheless, computing the global gradient in (1) requires noiseless aggregation of all the local gradients, i.e., they need to be perfectly received at the PS, and aggregated with uniform weights $\frac{1}{N}$. However, in practice, local gradients are obtained through a noisy wireless channel, introducing errors in the estimation of the global gradient $\overline{\boldsymbol{g}}_{t}$, as discussed next.
@@ II-A Over-the-air transmission over a fading MAC
We model the channel from device $m$ to the PS in round $t$ as flat Rayleigh fading, $h_{m,t}\sim\mathcal{CN}(0,\Lambda_{m})$, independent across devices and i.i.d. over $t$. Here, the parameter $\Lambda_{m}$ represents the average channel gain, which is dependent on large-scale propagation conditions and is assumed to be constant throughout the FL runtime. We assume the PS has knowledge of this statistical CSI, i.e., $\{\Lambda_{m}\}$, but not the instantaneous CSI $\{h_{m,t}\}$. Critically, unlike existing works [3, 21, 23, 16] assuming identical average path loss across devices ($\Lambda_{m}{=}\Lambda_{n},\forall m,n\in[N]$}), we consider a heterogeneous wireless environment where devices experience distinct path losses and study uplink transmission of local gradients via over-the-air (OTA) aggregation. OTA-FL exploits the superposition property of the wireless multiple-access channel to perform joint computation-and-communication, enabling single-shot aggregation at the PS. Following [21, 3, 24, 23, 14, 16], we assume devices are perfectly synchronized during gradient upload. Denoting by $\mathbf{x}_{m,t}$ the signal transmitted by device $m$ in round $t$, the PS receives the signal
@@
Here, $E_{s}$ is the maximum per-sample energy budget, and $G_{\text{max}}$ is an upper bound on $\|\boldsymbol{g}_{m,t}\|$ (see Assumption 2). By construction, a device remains silent when $|h_{m,t}|<\frac{G_{\text{max}}\gamma_{m}}{\sqrt{dE_{s}}}$ to ensure the energy constraint $\|\mathbf{x}_{m,t}\|^{2}/d\leq E_{s}\,,\forall m,t$, where $d$ is the model dimension. The decision in (5) is decentralized using the instantaneous uplink CSI $h_{m,t}$, which can be acquired with negligible overhead via a downlink pilot, by leveraging channel reciprocity [19]. With this design, the PS estimates the global gradient (1) as
@@ II-B Biased Over-the-Air-FL
in place of the ideal SGD update in (2). Conditioned on $\mathbf{w}_{t}$ and the mini-batch selections, the received signal satisfies $\mathbb{E}[\mathbf{y}_{t}]=\sum_{m\in[N]}\alpha_{m}\,\boldsymbol{g}_{m,t}$ with $\alpha_{m}\triangleq\gamma_{m}\,\mathbb{E}[\chi_{m,t}]={\gamma_{m}\exp\{-\tfrac{\gamma_{m}^{2}G_{\text{max}}^{2}}{d\,\Lambda_{m}\,E_{s}}\}}$, where the expectation is over channel fading and receiver noise. Choosing the post-scaler $\alpha\triangleq\sum_{m\in[N]}\alpha_{m}$ yields a global gradient estimate such that the expected gradient estimate is a convex combination of local stochastic gradients,
$\tilde{\boldsymbol{g}}_{t}\;\triangleq\;\mathbb{E}[\hat{\boldsymbol{g}}_{t}\mid\mathbf{w}_{t},\{\mathcal{B}_{m,t}\}_{m}]\;=\;\sum_{m\in[N]}p_{m}\,\boldsymbol{g}_{m,t},$ (8)
with weights $p_{m}\triangleq\alpha_{m}/\alpha$, $0\leq p_{m}\leq 1$, and $\sum_{m}p_{m}=1$. We interpret $p_{m}$ as the average participation level of device $m$ induced by the OTA-FL design. Consequently, the update rule in (7) acts as an SGD method that, on average, descends along $\tilde{\boldsymbol{g}}_{t}$ in place of the ideal aggregate $\overline{\boldsymbol{g}}_{t}$ from (1). Therefore, these updates minimize a different objective function than (P), on average, given by
@@
It is to be highlighted that prior OTA-FL works [22, 17, 3, 21, 23, 16] either assume homogeneous large-scale wireless conditions or enforce zero-bias aggregation, thereby ensuring uniform participation $p_{m}=\frac{1}{N}$ for all $m\in[N]$, so that minimizing ($\tilde{\text{P}}$) becomes equivalent to (P). While effective under homogeneous conditions, these designs become constrained by the weakest channel device in heterogeneous wireless networks, leading to high-variance model updates (see, e.g., [24, 1, 6]). Moreover, under wireless heterogeneity, such schemes effectively minimize a different objective, leading to objective inconsistency [18] and a non-negligible model bias. As a result, their convergence guarantees do not directly apply to the scenario considered in this paper.
Our proposed OTA-FL aggregation design generalizes existing schemes by introducing a controllable, non-zero average bias, thereby subsuming zero-bias designs ($p_{m}=\tfrac{1}{N}$) as a special case. This formulation exposes an explicit bias–variance trade-off that can be optimized to improve convergence. While some recent works [6, 25] also consider biased OTA-FL, they typically allow a generic, unstructured bias that is difficult to analyze and control. In contrast, building on our prior work [1], we adopt a structured, time-invariant bias that admits tractable finite-time guarantees. Note that, unlike our prior study [1], which treated (strongly) convex objectives and offered heuristic designs, here we derive a convergence bound for non-convex objectives and optimize the bias-variance trade-off via an efficient SCA method, which is done in the next section.